consider a neutral soherically conducting shell of inner radius r1 and outer radius r2. a point charge +q, is placed at tge center (r=0) of the spherically conducting shell. Answer the following questions symbollically in terms of k,r1,r2, and q. a) what is the electric field for r>r1 b) what is the electric field for r2>r>r1? c) what is the electric field for r>r2? d) what is the surface charge density omega1, on the inner surface of the shell? e) what is the surface charge density omega2, on the outside of the shell? f) what is the net charge on the spherical conducting shell?
Answer #1

A) for r>r1,

E= kq/r^2

B) for r1<r<r2 electric field will be zero as no charge resides in this area if we apply guass law.

C) for r> r2

Electric field E = kq/r^2

D) as charge induced on inner surface will be -q so

omega 1 = -q/4(pi)r1^2

Here pi= 3.14

E) charge induced on outer surface is +q

So

Omega2 = q/4(pi)r2^2

Here r2^2 means r2×r2 ie square of r2

F) net charge on spherical shell is +q as the charge at center and charge on inner surface will sum to zero